OpenStudy (anonymous):

Find f F"(x)=8x^3+5. F(1)=0. F'(1)=8

OpenStudy (anonymous):

F"(x)=8x^3+5. F'(x) = int (8x^3+5) dx F'(x) = (8x^4)/4+5x + C F'(x) = 2x^4 + 5x + C because of F'(1) = 8, then F'(x) = 2x^4 + 5x + C F'(1) = 2(1)^4 + 5(1) + C = 8 2(1)^4 + 5(1) + C = 8 2 + 5 + C = 8 C = 8-2-5 C = 8-7 C = 1

OpenStudy (anonymous):

Thank you I get it now

OpenStudy (anonymous):

So F'(x) = 2x^4 + 5x + C F'(x) = 2x^4 + 5x + 1 now, we'll find F(X) F(x) = int (F'(x)) dx F(x) = int ( 2x^4 + 5x + 1) dx F(x) = (2x^5)/5 + (5x^2)/2 + x + C F(x) = (2x^5)/5 + (5x^2)/2 + x + C because of F(1) = 0, then F(x) = (2x^5)/5 + (5x^2)/2 + x + C F(1) = (2(1)^5)/5 + (5(1)^2)/2 + (1) + C = 0 (2(1)^5)/5 + (5(1)^2)/2 + (1) + C = 0 (2/5) + (5/2) + 1 + C = 0 C = (2/5) - (5/2) - 1 C = -3.1 So F(x) = (2x^5)/5 + (5x^2)/2 + x + C F(x) = (2x^5)/5 + (5x^2)/2 + x -3.1

OpenStudy (anonymous):

you are welcome